# Zero modes and index theorems for non-Hermitian Dirac fermions

### v1

@@ V Summary and discussion
To summarize, I here extend the jurisdiction of three prominent index theorems for quasi-relativistic Dirac fermions to open or NH systems, for which the NH Dirac operators besides featuring the Lorentz invariance also display a purely real energy eigenvalue spectrum over an extended NH parameter regime. Such NH planar Dirac systems, when placed in external magnetic fields continue to host robust zero-energy modes, the number of which is equal to the magnetic flux quanta enclosed by the system, irrespective of the strength of the NH perturbation in the system. On the other hand, when massive NH Dirac fermions sustain a topologically non-trivial texture of the mass order in the real space, such as a domain-wall in one dimension and vortices in two dimensions, the system still honors the Jackiw-Rebbi and Jackiw-Rossi index theorems, respectively, as in the Hermitian systems, but only when the effective Fermi velocity for NH Dirac fermions is purely real.
@@
These findings should be of far reaching consequences in open or NH quasi-relativistic systems. For example, existence of zero-energy modes in the presence of external magnetic fields of arbitrary shape strongly suggests that the magnetic catalysis mechanism for the nucleation of chiral symmetry breaking mass orders should remain operative in weakly interacting NH planar Dirac systems [20, 53, 54]. Dynamic symmetry breaking following this mechanism causes insulation in the system at half-filling and gives rise to the notion of competing orders within the zero-energy manifold. Zero-energy bound states, localized near the core of topological defects of mass order, can give rise to charge fractionalization (for charged Dirac fermions) and localized Majorana modes (for neutral Bogoliubov Dirac quasiparticles) in open quasi-relativistic systems. The findings related to the Jackiw-Rossi index theorem extend directly to the situation with a line vortex in three-dimensional NH massive Dirac and Weyl systems [50, 55], suggesting the presence of dispersive plane-wave modes along the vortex line that are localized around the vortex core as long as the effective Fermi velocity of NH Dirac fermions is real. In this case, the celebrated Callan-Harvey mechanism becomes operative due to the presence of one-dimensional modes along the vortex core [56], leading to a flow of nondissipative quantized current, given by $j_{z}=ne^{2}E_{z}/(2\pi)$, when an electric field ($E_{z}$) is applied in the $z$ direction, where $n$ is the number of zero-energy localized modes bound to vortex core. Such a current will be supplied radially into the vortex core which is captured by axion electrodynamics term. A detailed derivation of this phenomenon is, however, left for a future investigation.
@@
The simplicity of the construction of the NH Dirac operator, see Sec. I.1, makes the theoretical predictions from this work testable in experiments. For example, a collection of Hermitian massless Dirac fermions can be realized on graphene’s honeycomb lattice, which can be captured from the nearest-neighbor tight-binding Hamiltonian. In a two-component spinor basis $\Psi^{\top}_{\bf k}=(c_{A},c_{B})({\bf k})$, where $c_{A}({\bf k})$ and $c_{B}({\bf k})$ are the fermionic annihilation operators on the sites of two interpenetrating triangular sublattices $A$ and $B$, respectively, of the honeycomb lattice such a tight-binding model with hopping amplitude $t_{0}$ leads to the following Bloch Hamiltonian
@@
where $f({\bf k})=\exp[i{\bf k}\cdot{\bf b}_{1}]+\exp[i{\bf k}\cdot{\bf b}_{2}]+\exp[i{\bf k}\cdot{\bf b}_{3}]$, and ${\bf b}_{1}=(1/\sqrt{3},1)a/2$, ${\bf b}_{2}=(1/\sqrt{3},-1)a/2$, and ${\bf b}_{3}=(-1/\sqrt{3},0)a$ are the nearest-neighbor vectors. Here $a$ is the lattice spacing and $\star$’ denotes the complex conjugation. The above Bloch Hamiltonian fosters massless Dirac fermions around six corners of the hexagonal Brillouin zone. With a specific choice of $M={\rm diag.}(1,-1)$ (yielding a staggered potential between two sublattices of the honeycomb lattice [46]) in Eq. (2), I then realize a collection of NH massless Dirac fermions, where the non-Hermiticity results from the imbalance between the hopping amplitudes in the opposite directions between two sublattices. Such a platform yields the ideal setup where NH generalization of the Aharonov-Casher index theorem can be tested at least on optical honeycomb lattices [57] on which the orbital effects of magnetic field can be engineered by synthetic gauge fields [58, 59] and hopping imbalance yielding non-Hermiticity [60, 61] can in principle be emulated.
@@
In the same setup the Haldane’s quantum anomalous Hall insulator is represented by a mass matrix $M={\rm diag.}(1,-1)g(\bm{k})$, where $g(\bm{k})=\exp[i{\bf k}\cdot{\bf a}_{1}]+\exp[i{\bf k}\cdot{\bf a}_{2}]+\exp[i{\bf k}\cdot({\bf a}_{1}+{\bf a}_{2})]$ and ${\bf a}_{1}$ and ${\bf a}_{2}$ are Bravais lattice vectors of the triangular lattice [47]. With such a choice of $M$, the corresponding NH Dirac operator from Eq. (2) also yields an imbalance of the hopping amplitudes between the nearest-neighbor sites in the opposite directions. Since Haldane’s anomalous Hall insulator order has been engineered on optical honeycomb lattice [62], it is natural to expect that such an intra-sublattice circulating current pattern can be utilized to engineer the corresponding NH Dirac operator therein following the general protocol from Sec. I.1. On such a setup, a Y-junction of Kekulé valence bond order (yielding lattice regularization of a U(1) mass vortex) can bind localized zero-energy modes to feature charge fractionalization for NH Dirac fermions. Besides optical honeycomb lattices, artificial honeycomb lattices also constitute another promising platform where the predicted zero-modes bound to mass vortex causing charge fractionalization can be observed since on such systems Kekulé valence bond order has already been realized [63, 64]. Altogether, the Lorentz invariant NH Dirac theory constitute an ideal platform where the footprints of various celebrated index theorems in open quasi-relativistic systems can be tested theoretically and their experimental verification should be within the reach of currently accessible facilities to engineer desired NH systems.
@@ Acknowledgements.
B.R. was supported by NSF CAREER Grant No. DMR-2238679 (USA) and thanks ANRF, India, for support through the Vajra scheme VJR/2022/000022. The author is thankful to Christopher A. Leong and Vladimir Juričić for critical reading of the manuscript.

### v2

@@ V Summary and discussion
To summarize, I here extend the jurisdiction of three prominent index theorems for quasirelativistic Dirac fermions to open or NH systems, for which the NH Dirac operators besides featuring the Lorentz invariance also display a purely real energy eigenvalue spectrum over an extended NH parameter regime. Such NH planar Dirac systems, when placed in external magnetic fields continue to host robust zero-energy modes, the number of which is equal to the number of magnetic flux quanta enclosed by the system, irrespective of the strength of the NH perturbation and the profile of the magnetic field in the system. On the other hand, when massive NH Dirac fermions sustain a topologically nontrivial texture of the mass order in the real space, such as a domain wall in one dimension and vortices in two dimensions, the system still honors the Jackiw-Rebbi and Jackiw-Rossi index theorems, respectively, as in the Hermitian systems, but only when the effective Fermi velocity for NH Dirac fermions is purely real. In these cases, on the other hand, when the Fermi velocity becomes imaginary, the zero-energy modes become plane waves, a phenomenon exclusively occurring in NH systems. In future, it will be worthwhile finding appropriate interacting models that can sustain the desired “commuting class mass” orders in NH Dirac system in which one can then study the imprints of topological defects [38, 39, 40, 41, 42, 57].
I also notice that for $|\alpha|<1$ and even more interestingly when $|\alpha|>1$ (yielding an imaginary Fermi velocity), the vortex core with at least $|n|=1$ continues to support a single normalizable zero-energy mode when the system is subject to regular magnetic fields (not the deconfining one) of arbitrary spatial profile, the effects of which are incorporated by taking $i\partial_{j}-A_{j}$ for $j=x$ and $y$ (minimal coupling with the magnetic vector potential $\bm{A}$) in Eq. (IV). The resulting NH gauged-Dirac operator enjoys the unitary particle-hole symmetry $\{H^{\rm NH}_{\rm Dir}(\bm{k}\to-i{\bm{\nabla}}-\bm{A},\alpha,n),\Gamma_{5}\}=0$, which protects the zero modes of the schematic form shown in Eq. (32). This conclusion follows from the one in Hermitian systems [58, 59, 60], with the solutions expressed in terms of parabolic cylindrical functions [61], when the magnetic field is uniform. The outcomes can be justified in the following way. Notice that in the presence of magnetic fields ($\bm{B}$) in a system with a mass vortex, there are two length scales, namely the magnetic length ($\ell_{B}$) and a length scale due to the mass order $\ell_{\Delta}=(\Delta_{0}/\sqrt{1-\alpha^{2}})^{-1}$ that together determine the localization of the zero-energy mode bound to such defect. For $|\alpha|<1$, both $\ell_{B}$ and $\ell_{\Delta}$ are real, and the solutions of the zero-energy modes in NH systems are smoothly connected to the one in Hermitian setups. For $|\alpha|>1$, only $\ell_{B}$ is real, which then solely determines the localization of the normalizable zero-energy modes. Only for the sake of brevity do I not show here the explicit solutions of such a mode.
@@
These findings should be of far reaching consequences in open or NH quasirelativistic systems. For example, existence of zero-energy modes in the presence of external magnetic fields of arbitrary shape strongly suggests that the magnetic catalysis mechanism for the nucleation of chiral symmetry breaking mass orders should remain operative in weakly interacting NH planar Dirac systems [21, 62, 63]. Dynamic symmetry breaking following this mechanism causes insulation in the system at half-filling and gives rise to the notion of competing orders within the zero-energy manifold. Zero-energy bound states, localized near the core of topological defects of mass orders, can give rise to charge fractionalization (for charged Dirac fermions) and localized Majorana modes (for neutral Bogoliubov Dirac quasiparticles) in open quasirelativistic systems. The localized zero-energy modes near the core of topological defects of mass orders can also foster additional and distinct local mass order parameters [34, 35, 36]. Therefore, when such defects proliferate in the system, local mass orders can acquire a global coherence, leading the notion of deconfined quantum phase transition, which thus far has been discussed exclusively in closed or Hermitian systems [64, 65, 66, 67]. The present discussion, therefore, opens an uncharted territory of such continuous deconfined phase transition between two distinct ordered phases via the proliferation of real space topological defects in open or NH systems. The index theorems and the existence of zero-energy modes in NH systems are also germane in the context of topological phases of matter [52].
The findings related to the Jackiw-Rossi index theorem extend directly to the situation with a $z$-directional line vortex in three-dimensional NH massive Dirac and Weyl systems [54, 68], suggesting the presence of dispersive plane-wave modes along the vortex line that are localized around the vortex core as long as the effective Fermi velocity of NH Dirac fermions is real. In this case, the celebrated Callan-Harvey mechanism becomes operative due to the presence of one-dimensional modes along the vortex core [69], leading to a flow of nondissipative quantized current, given by $j_{z}=ne^{2}E_{z}/(2\pi)$, when an electric field ($E_{z}$) is applied in the $z$ direction, where $n$ is the number of zero-energy localized modes bound to the vortex core. Such a current will be supplied radially into the vortex core which can be captured by an axion electrodynamics term. A detailed derivation of this phenomenon is, however, left for a future investigation. The present discussion can also be extended to string-type configurations for a general fermionic mass matrix for neutrinos in an extended standard model [70], but in a NH formalism.
@@
The simplicity of the construction of the NH Dirac operator, see Sec. I.1, makes the theoretical predictions from this work testable in tabletop experiments. For example, a collection of Hermitian massless Dirac fermions can be realized on graphene’s honeycomb lattice, which can be captured from the nearest-neighbor tight-binding Hamiltonian. In a two-component spinor basis $\Psi^{\top}_{\bf k}=(c_{A},c_{B})({\bf k})$, where $c_{A}({\bf k})$ and $c_{B}({\bf k})$ are the fermionic annihilation operators on the sites of two interpenetrating triangular sublattices $A$ and $B$, respectively, of the honeycomb lattice such a tight-binding model with hopping amplitude $t_{0}$ leads to the following Bloch Hamiltonian
@@
where $f({\bf k})=\exp[i{\bf k}\cdot{\bf b}_{1}]+\exp[i{\bf k}\cdot{\bf b}_{2}]+\exp[i{\bf k}\cdot{\bf b}_{3}]$, and ${\bf b}_{1}=(1/\sqrt{3},1)a/2$, ${\bf b}_{2}=(1/\sqrt{3},-1)a/2$, and ${\bf b}_{3}=(-1/\sqrt{3},0)a$ are the nearest-neighbor vectors. Here, $a$ is the lattice spacing and $\star$” denotes the complex conjugation. The above Bloch Hamiltonian fosters massless Dirac fermions around six corners of the hexagonal Brillouin zone. With a specific choice of $M={\rm diag.}(1,-1)$ (yielding a staggered potential between two sublattices of the honeycomb lattice [50]) in Eq. (2), I then realize a collection of NH massless Dirac fermions, where the non-Hermiticity results from the imbalance between the hopping amplitudes in the opposite directions between two sublattices. Such a platform yields the ideal setup where NH generalization of the Aharonov-Casher index theorem can be tested at least on optical honeycomb lattices [71] on which the orbital effects of magnetic field can be engineered by synthetic gauge fields [72, 73] and hopping imbalance yielding non-Hermiticity [74, 75] can in principle be emulated.
@@
In the same setup the Haldane’s quantum anomalous Hall insulator is represented by a mass matrix $M={\rm diag.}(1,-1)g(\bm{k})$, where $g(\bm{k})=\sin(i{\bf k}\cdot{\bf a}_{1})+\sin(i{\bf k}\cdot{\bf a}_{2})+\sin(i{\bf k}\cdot[{\bf a}_{2}-{\bf a}_{1}])$ and ${\bf a}_{1}=(\sqrt{3},-1)a/2$ and ${\bf a}_{2}=(0,1)a$ are Bravais lattice vectors of the triangular lattice [51]. With such a choice of $M$, the corresponding NH Dirac operator from Eq. (2) also yields an imbalance of the hopping amplitudes between the nearest-neighbor sites in the opposite directions. Since Haldane’s anomalous Hall insulator order has already been engineered on optical honeycomb lattice [76], it is natural to expect that such an intrasublattice circulating current pattern can be utilized to engineer the corresponding NH Dirac operator therein following the general protocol from Sec. I.1. On such a setup, a Y junction of Kekulé valence bond order [yielding lattice regularization of a U(1) mass vortex] can bind localized zero-energy modes to feature charge fractionalization for NH Dirac fermions. Furthermore, three-dimensional Weyl fermions have also been realized on optical lattices recently [77], constituting the ideal stage on which an extension of the Callan-Harvey mechanism, for example, in a NH Weyl system can be showcased, as conjectured from the present discussion, especially given the recent success in engineering non-Hermiticity in cold atomic systems [75].
Besides optical honeycomb lattices, artificial honeycomb lattices also constitute another promising platform where the predicted zero modes bound to a mass vortex causing charge fractionalization can be observed since on such systems the Kekulé valence bond order has already been realized [78, 79]. Altogether, the Lorentz invariant NH Dirac theory constitutes an ideal platform where the footprints of various celebrated index theorems in open quasirelativistic systems can be tested theoretically and their tabletop experimental verification should be within the reach of currently accessible facilities to engineer desired NH quasirelativistic lattices with balanced gain and loss (required for all-real eigenvalues of a NH operator).
Extending the index theorems to non-Lorentz invariant NH systems should be a interesting generalization of the current pursuit. Note that Lorentz symmetry can be violated in numerous ways and encompassing all of them certainly goes beyond the scope of the present discussion. Nonetheless, here I discuss one such case in two dimensions ($d=2$) where the conclusions from the present work can be extended straightforwardly to predict the existence of topologically robust zero-energy states, following the existing results in Hermitian systems and the general principle of constructing the corresponding NH operator from Sec. I.1. For example, one can consider the following generalization of the Dirac Hamiltonian from Eq. (1) describing an $s$th order touching of the valence and the conduction bands for which the Hamiltonian reads as
$H_{s}(\bm{k})=\Lambda_{s}\sum_{j=1,2}\Gamma_{j}\;d^{s}_{j}(\bm{k})$ (53)
with $\bm{d}^{s}(\bm{k})=(d^{s}_{1},d^{s}_{2})(\bm{k})=|k|^{s}(\cos(s\varphi_{\bm{k}}),\sin(s\varphi_{\bm{k}}))$, where $\varphi_{\bm{k}}=\tan^{-1}(k_{y}/k_{x})$ and $\Lambda_{s}$ bears the appropriate dimension such that $H_{s}$ carries the dimension of energy. For example, when $s=2$, the parameter $\Lambda_{s}$ has the dimension of inverse mass. Throughout this work, I considered only Lorentz symmetric systems with $s=1$ for which $\Lambda_{s}\equiv v_{{}_{\rm H}}$. Due to the nonlinear energy momentum relation $E(\bm{k})=\pm\Lambda_{s}\;|\bm{k}|^{s}$, the Lorentz symmetry gets violated for any integer $s>1$. Nonetheless, the entire construction of the corresponding NH operator now follows straightforwardly from Sec. I.1 by taking $H_{\rm Dir}(\bm{k})\to H_{s}(\bm{k})$ therein, as there exist a plethora of mass order matrices $M$ that satisfy $\{H_{s}(\bm{k}),M\}=0$.
The Aharonov-Cashed index theorem in such a Hermitian system guarantees the existence of $sN$ number of topologically robust zero-energy modes whose spatial localization is solely determined by the profile of the external (uniform or inhomogeneous) magnetic fields ($\Lambda_{s}$ playing no role) [20], where $N$ is the number of magnetic flux quanta enclosed by the system. As the NH counterpart features the same unitary particle-hole symmetry (generated by $\sigma_{3}$), I expect the corresponding operator to support $sN$ number of normalizable zero-energy modes, irrespective of the strength of the non-Hermiticity $\alpha$, as shown explicitly for $s=1$ here. The Jackiw-Rossi index theorem in such Hermitian models with an underlying mass vortex of vorticity one yields $s$ number of normalizable zero-energy modes [80, 81], which are expected to survive (due to the unitary particle-hole symmetry, generated by $\Gamma_{5}$) in NH systems as long as $|\alpha|<1$. Existence of such $s$ number of zero-energy modes can be responsible for charge $se$ superconductivity [82], which should be also operative in the analogous NH setups. Explicit demonstrations of these outcomes are, however, left for future investigations. For other index theorems in Lorentz-violating NH systems, see Ref. [83].
@@ Acknowledgements.
The author is thankful to Christopher A. Leong and Vladimir Juričić for critical reading of the manuscript. B.R. was supported by NSF CAREER Grant No. DMR-2238679 (USA) and thanks ANRF, India, for support through the Vajra scheme VJR/2022/000022.
Data Availability
No data were created or analyzed in this study.